# SA-MLP: Enhancing Point Cloud Classification with Efficient Addition and Shift Operations in MLP Architectures

### v1

@@ 3.4 Combined Model: ShiftAdd-MLP (SA-MLP)
A critical distinction between SA-MLP and ShiftAddNet [12] lies in the treatment and optimization of the shift and adder layers. In ShiftAddNet [12], the shift layer is frozen during training, preventing it from contributing to learning and reducing it to a source of random perturbation, which can degrade overall network performance. In contrast, SA-MLP allows both shift and adder layers to actively participate in the learning process by using differentiated optimization strategies tailored to the characteristics of each layer type. Although these strategies are only briefly mentioned here, they are vital to SA-MLP’s superior performance and will be detailed in the following subsection (Sec. 3.5), which discusses the optimization techniques for Shift-MLP, Add-MLP, and SA-MLP.
@@
3.5 Optimization Strategies for Shift and Adder Layers in SA-MLP
@@
3.5.1 Shift Layer Optimization
@@
3.5.2 Adder Layer Optimization
@@ 4.1 Training Configuration
4.2 Classification Task
In this subsection, we evaluate the performance of the SA-MLP model proposed in this study against several state-of-the-art methods on the ModelNet40 classification task. The results of this comparative analysis are summarized in Tab. 1.
@@ 4.2 Classification Task
Table 1: Results for the ModelNet40 classification task.
Tab. 1 illustrates the performance of various MLP-based methods, including PointNet [7], PointNet++[13], and PointNext[30]. Remarkably, the SA-MLP model proposed in this study surpasses these established methods, achieving a classification accuracy of 93.9%. This notable improvement underscores the efficacy of combining both shift and add operations within the MLP framework, as demonstrated in this research. The inclusion of these operations enhances the model’s ability to extract and process point cloud features, contributing to its superior performance.
@@
4.3 Performance Across Point Cloud Densities
The table 2 presents the classification accuracy of four models—Mul-MLP, Shift-MLP, Add-MLP, and SA-MLP—on the ModelNet40 dataset, evaluated with varying point cloud densities: 1024, 512, 256, and 128 points. Across all models, there is a clear trend: as the number of points decreases, the classification accuracy generally declines. This pattern is expected, as a lower number of points reduces the amount of information available for feature extraction, making it more challenging for the models to accurately classify the point clouds. However, the extent of this decline varies among the models, providing insight into their robustness to changes in point cloud density.
@@ 4.3 Performance Across Point Cloud Densities
Table 2: Classification accuracy of Mul-MLP, Shift-MLP, Add-MLP, and SA-MLP models on ModelNet40 across varying point cloud densities.
@@
4.4 Analysis of Gradient Regularization Across Layer Types
This subsection analyzes the gradient regularization applied across different layer types, specifically focusing on the adder layers compared to multiplication-based and shift-based layers. The analysis examines the root-mean-square (RMS) values of the gradients across various embedding and encoder layers, as shown in Tab. 3. The table categorizes the different layers based on their operational types: multiplication (mul.), shift, and addition (add), with special consideration for the adder layers’ vanilla gradients before adaptive modulation (add (van.)).
@@ 4.4 Analysis of Gradient Regularization Across Layer Types
Table 3: Comparison of Root-Mean-Square (RMS) values of gradients across embedding and encoder layers for Mul-MLP, Shift-MLP, Add-MLP, and SA-MLP models.
@@
4.5 Visualization Analysis of Encoder Features
@@
4.6 Weight Distribution Analysis Across Different Layer Types
@@
5 Conclusion
This study proposed a series of multiplication-free MLP-based models targeted at improving the computational efficiency of point cloud classification tasks by replacing traditional multiplication operations with shift and add operations. Through extensive experimentation, we demonstrated that while individual models like Shift-MLP and Add-MLP offered comparable accuracy to multiplication-based models, the hybrid SA-MLP model, which strategically integrates both shift and add operations, emerged as the best performer. It not only outperformed the baseline Mul-MLP but also surpassed several state-of-the-art methods on the ModelNet40 classification benchmark, affirming the potential of multiplication-free architectures in high-accuracy, resource-efficient deep learning applications.
However, despite the promising results, the study also encountered significant limitations that warrant discussion. The training of the adder layer was notably slower compared to traditional multiplication-based models. This slowdown can be attributed to the current limitations in deep learning infrastructure, particularly with regard to the lack of optimization for non-standard operations like addition-based networks. Modern deep learning libraries, such as CUDA and CuDNN, have been extensively optimized for traditional convolutional operations through dedicated hardware acceleration, advanced algorithms, and auto-tuning capabilities, which significantly boost the performance of CNNs. In contrast, the addition and absolute value operations central to adder layers do not benefit from these optimizations, leading to inefficiencies during both forward inference and backpropagation. Although CUDA programming was employed in this study to optimize these operations, the results indicate that the training and inference speeds on GPU platforms are still inferior to those of traditional matrix multiplication-based models.
These limitations highlight the challenges of generalizing addition-based networks to more complex point cloud analysis tasks, such as semantic segmentation. While the SA-MLP model presents a viable solution for point cloud classification, its broader applicability remains constrained by the current state of deep learning infrastructure. Future research should explore the development of more optimized frameworks and hardware support for non-standard operations, which could unlock the full potential of addition-based networks. Moreover, investigating alternative strategies to mitigate the computational overhead of adder layers could pave the way for their integration into more complex and computationally intensive tasks. Despite these challenges, this study provides a foundation for further exploration into efficient, multiplication-free neural network architectures, with the potential to significantly impact the field of deep learning in resource-constrained environments.

### v2

@@ III-D Combined Model: ShiftAdd-MLP (SA-MLP)
A critical distinction between SA-MLP and ShiftAddNet [13] lies in the treatment and optimization of the shift and adder layers. In ShiftAddNet [13], the shift layer is frozen during training, preventing it from contributing to learning and reducing it to a source of random perturbation, which can degrade overall network performance. In contrast, SA-MLP allows both shift and adder layers to actively participate in the learning process by using differentiated optimization strategies tailored to the characteristics of each layer type. Although these strategies are only briefly mentioned here, they are vital to SA-MLP’s superior performance and will be detailed in the following subsection (Sec. III-E), which discusses the optimization techniques for Shift-MLP, Add-MLP, and SA-MLP.
@@
III-E Optimization Strategies for Shift and Adder Layers in SA-MLP
@@
III-E1 Shift Layer Optimization
@@
III-E2 Adder Layer Optimization
@@ IV-A Training Configuration
IV-B Classification Task
In this subsection, we evaluate the performance of the SA-MLP model proposed in this study against several state-of-the-art methods on the ModelNet40 classification task. The results of this comparative analysis are summarized in Tab. I.
@@ IV-B Classification Task
TABLE I: Results for the ModelNet40 classification task.
Tab. I illustrates the performance of various MLP-based methods, including PointNet [7], PointNet++[14], and PointNext[28]. Remarkably, the SA-MLP model proposed in this study surpasses these established methods, achieving a classification accuracy of 93.9%. This notable improvement underscores the efficacy of combining both shift and add operations within the MLP framework, as demonstrated in this research. The inclusion of these operations enhances the model’s ability to extract and process point cloud features, contributing to its superior performance.
@@
IV-C Performance Across Point Cloud Densities
The table II presents the classification accuracy of four models—Mul-MLP, Shift-MLP, Add-MLP, and SA-MLP—on the ModelNet40 dataset, evaluated with varying point cloud densities: 1024, 512, 256, and 128 points. Across all models, there is a clear trend: as the number of points decreases, the classification accuracy generally declines. This pattern is expected, as a lower number of points reduces the amount of information available for feature extraction, making it more challenging for the models to accurately classify the point clouds. However, the extent of this decline varies among the models, providing insight into their robustness to changes in point cloud density.
@@ IV-C Performance Across Point Cloud Densities
TABLE II: Classification accuracy of Mul-MLP, Shift-MLP, Add-MLP, and SA-MLP models on ModelNet40 across varying point cloud densities.
@@
IV-D Analysis of Gradient Regularization Across Layer Types
This subsection analyzes the gradient regularization applied across different layer types, specifically focusing on the adder layers compared to multiplication-based and shift-based layers. The analysis examines the root-mean-square (RMS) values of the gradients across various embedding and encoder layers, as shown in Tab. III. The table categorizes the different layers based on their operational types: multiplication (mul.), shift, and addition (add), with special consideration for the adder layers’ vanilla gradients before adaptive modulation (add (van.)).
@@ IV-D Analysis of Gradient Regularization Across Layer Types
TABLE III: Comparison of Root-Mean-Square (RMS) values of gradients across embedding and encoder layers for Mul-MLP, Shift-MLP, Add-MLP, and SA-MLP models.
@@
IV-E Visualization Analysis of Encoder Features
@@
IV-F Weight Distribution Analysis Across Different Layer Types
@@
V Conclusion
This study presented a series of multiplication-free MLP-based models tailored to the classification of 3D point clouds, with a primary focus on enhancing computational efficiency and reducing energy consumption, particularly in resource-constrained environments such as sensor systems and edge devices. By replacing traditional multiplication operations with shift and add operations, the proposed models, including Shift-MLP, Add-MLP, and the hybrid SA-MLP, demonstrated comparable or superior classification performance compared to conventional models. Among these, SA-MLP, which combines both shift and add operations, outperformed the baseline Mul-MLP and several state-of-the-art methods on the ModelNet40 benchmark, illustrating the potential of multiplication-free architectures as viable solutions for point cloud classification tasks, especially in sensor systems where computational resources are often constrained.
Despite these promising results, the study also identified several limitations. The training of adder layers was slower compared to traditional multiplication-based models, primarily due to the lack of optimization for addition and shift operations within current deep learning frameworks. While existing libraries such as CUDA and CuDNN are highly optimized for matrix multiplication and convolutional operations, non-standard operations like addition-based computations still lack comparable support, leading to inefficiencies during both training and inference on GPU platforms. These challenges highlight the gap in existing infrastructure for integrating multiplication-free models into widely-used deep learning workflows, thus requiring further development to fully harness the potential of these architectures in real-world applications.
To address these limitations, future research should focus on optimizing deep learning frameworks and hardware accelerators specifically for non-standard operations, such as addition and shift-based computations. Tailored optimizations for these operations could substantially improve the performance and scalability of multiplication-free architectures, making them more practical for real-time sensor data processing and enabling their broader adoption in resource-constrained environments. Furthermore, exploring advanced techniques to reduce the computational overhead of the adder layers could make these models more feasible for large-scale point cloud processing tasks, including semantic segmentation and large-scale 3D point cloud classification.
In conclusion, this work represents a significant step toward the integration of multiplication-free architectures into point cloud analysis, demonstrating that such models can substantially reduce computational consumption while maintaining high accuracy. Although the current study is preliminary, it provides a promising solution for efficient 3D point cloud classification, particularly in sensor-driven applications such as autonomous driving, robotics, and embedded systems. This work paves the way for future research to refine and optimize these models, with the potential to make a profound impact on point cloud analysis in resource-constrained environments, such as edge computing and sensor networks.